Mary decide to focus on selling luxury cars and needed to sell off one line of sports care: LaMancinis. She held a clearance sale. The first week of the sale, she and her associates sold half the LaMancinis. The second week, they again sold half the remaining sports cars on the lot and one more. Mary and the other sales reps then each choose one of the remaining cars to keep, and there were none left over. How many cars were on the lot when the sale began?
Answer by stanbon(75887)

Mary decide to focus on selling luxury cars and needed to sell off one line of sports care: LaMancinis. She held a clearance sale.
Assume they have "x" cars.
-------------------------------
The first week of the sale, she and her associates sold half the LaMancinis.
They now have "x/2" cars.
-------------------------------
The second week they again sold half the remaining sports cars on the lot and one more.
They now have (1/2)(x/2) -1 = (x/4) -1
--------------------------------------------
Mary and the other sales reps then each choose one of the remaining cars to keep, and there were none left over.
How many cars were on the lot when the sale began?
[(x/4)-1] = 2
(x/4) = 3
x = 12 cars
==================
Cheers,
Stan H.